# Colorful Minors

### v1

@@ 4.1 Colors in a surface
We should stress that the original formulation of proposition 4.1 is not stated in terms of -colorful graphs and -colorful boundary indices. However, a variant like proposition 4.1 is straightforward to obtain as in the original proof the boundary index is only used to assign a coloring to the non-vortex cells of a -rendition of which in this case would depend on Besides that, would not play any other role in its proof and can safely be disregarded thereafter.
@@ 4.2 Locally excluding a rainbow grid
Proposition 4.1 allows us to immediately deduce a strong local structure theorem for -colorful graphs excluding rainbow grids.
@@
We proceed by applying proposition 4.1 with and Notice that this means that This yields one of two outcomes as follows.
@@
Let be the minor model of and notice that is a truncation of We now apply theorem 3.1. If this results in a colorful minor model of a rainbow we have found every -colorful rainbow graph on vertices as a colorful minor in and thus, in particular, we have found the rainbow -grid. Hence, we may assume that we enter the second outcome of theorem 3.1 which yields a set of size at most
@@
4.3 The proof of Theorem 1.7
We are now ready for the proof of Theorem 1.7. The proof will largely follow a similar strategy as the one for theorem 1.5. Indeed, we will prove a stronger version similar to theorem 3.6 and its strengthened form theorem 3.7. Theorem 1.7 will then follow as a corollary. This slightly stronger version reads as follows:
@@
As announced, we strengthen theorem 4.3 further to allow for a smooth induction.
@@
First, we define the functions involved as follows where is the constant from proposition 2.3.
@@
We now proceed by induction on and start by discussing two fundamental cases. Exactly as in our proof of theorem 3.7, we first discuss the two fundamental cases where either has only few vertices, or These two cases are fully identical to the ones in theorem 3.7 and we therefore omit them here. Then two further cases follow:
Case 1, where has a small balanced separator, and Case 2, where does not have a small balanced separator. As before, for Case 1 there is no difference to the proof of theorem 3.7 and thus, we may omit it here.
@@
Since we may now apply proposition 2.3 to and obtain a -wall such that This further allows for the application of lemma 4.2 to the wall in In the following we distinguish three cases based on the three possible outcomes of lemma 4.2.
@@ 5 Excluding segregated grids
In this section, we are concerned with the proof of theorem 1.6, which is an analogue of the Grid Theorem of Robertson and Seymour for restrictive treewidth. Similar to the organization of section 4, we begin with some preliminary results and a “local” variant of the structure theorem before proceeding to the actual proof of theorem 1.6.
@@ 5.1 A homogeneous flat wall
This allows us to call upon proposition 2.4 to find either a model of controlled by in which case we would be done, or a set of size less than and an -subwall that is flat in
@@ 5.2 Locally excluding segregated grids
As it is custom, we begin by setting up our numbers. Let be the constant from proposition 2.3 and hfw be the function from lemma 5.1.
As a first step, we apply proposition 2.3 to and obtain a -wall Next, we apply lemma 5.1 to This either results in a -minor model with or in a set of size at most together with a homogeneous flat wall in of order
@@
We now apply theorem 3.1. This either yields a rainbow as a colorful minor in or a set of size at most such that the -big component of is restricted. With being a truncation of the second outcome matches the second outcome of our assertion. Hence we may assume to find a rainbow as a colorful minor. This, however, means that contains every -colorful graph on vertices as a colorful minor and this includes all -segregated grids. Hence, we are done with this case.
@@
Now let for some brick of Since is homogeneous, we know that also is homogeneous, so the choice of here is arbitrary. Let and for each let us denote by the set We now apply lemma 3.2 to the sets and the set asking for many paths for each member of This yields one of two outcomes. Either we find a linkage of order such that contains a --linkage of order for each and the paths in are all internalle vertex-disjoint from or there is a set of size less than such that there exists some and there is no --path in
@@
We claim that there exists an --linkage of size such that all internal vertices of the paths of belong to To see this, notice that in the absence of such a linkage proposition 2.1 would imply the existence of a separation of order at most between and However, each vertex of belongs to a horizontal path of that can be extended in at least different ways, intersecting only in the horizontal path, to a path that intersects all cycles at least of Hence, such a separator must leave at least one cycle of one way to get from some vertex of to and one way to get from to intact. Hence, deleting any set of size at most still leaves at least one --path intact and therefore, and cannot be separated by fewer than vertices in
Let be the perimeter of and be the -subwall of that avoids It is now easy to see that together with the paths from the paths from the horizontal paths of and the bricks of that intersect we may construct a -segregated grid as a colorful minor of See figure 4 for an illustration.
@@
5.3 A Grid Theorem for restrictive treewidth
We are now ready for the proof of theorem 1.6. Similar to our proofs of Theorems 3.6 and 1.7 we do this by proving a slightly different and stronger statement which, in turn, is implied by a further strengthening that lends itself better to induction.
@@ 5.3 A Grid Theorem for restrictive treewidth
contains a -segregated grid as a colorful minor and or
@@
Notice that the tree-decomposition from the second outcome of theorem 5.3 is a witness for Hence, theorem 5.3 does indeed act as an analogue of the Grid Theorem of Robertson and Seymour for restrictive treewidth.
We first prove that -segregated grids indeed act as lower bounds on the restrictive treewidth of a graph.
@@
For all positive integers and and every -colorful graph if contains a -segregated grid as a colorful minor, then
@@
All that is left to do is to prove the following strengthening of the upper bound part of theorem 5.3.
@@
We now proceed by induction on and start by discussing two fundamental cases. Exactly as in our proof of theorem 3.7, we first resolve the two fundamental cases where either has only few vertices, or in precisely the same way as before. Then two further cases follow:
Case 1, where has a small balanced separator, and Case 2, where does not have a small balanced separator. As before, for Case 1 there is no difference to the proof of theorem 3.7 and thus, we immediately move on to Case 2.
@@
In contrast to previous iterations of this proof, here we do not have to distinguish more subcases. We immediately may apply theorem 5.2 to and which either yields a -segregated grid as a colorful minor in – and there would be nothing more to prove – or it provides us with a set with such that the -big component of is segregated.
@@ 6 The Erdős-Pósa property for -colorful graphs
With our structural theorems in place, we are now ready to proceed with the proof of theorem 1.8. To this end, we first define, for every the sets and (section 6.1); we then introduce the class of crucial colorful graphs (section 6.1). The remainder of the section is devoted to proving that a -colorful graph has the Erdős-Pósa property if and only if it is crucial (theorem 6.19). In the proof, the set serves as the (colorful minor) obstruction set for the class of all -colorful graphs that satisfy the Erdős-Pósa property. The “upper bound” part of the proof is presented in section 6.3 (connected case) and section 6.4 (disconnected case), while the “lower bound” is established in section 6.5.
@@ 6.1 The sets and
We define by discarding all -colorful graphs with and every whose vertex set can be partitioned into two sets of size such that there are and For example, in the case where the only graph discarded is the one in the dashed square in figure 5.
@@
If then See figure 6 for an illustration.
@@
6.2 Crucial colorful graphs.
@@
Recalling Observations 6.2, 6.5, 6.8 and 6.11, one can see immediately that crucial colorful graphs are also closed under taking colorful minors.
@@
Taking into account lemma 6.15, it is easy to verify the following.
@@
By observation 6.16 we know that no member of contains another as a colorful minor. Moreover, by lemma 6.4 we have that no member of is color-facial, no member of is color-segmented by observation 6.7, no member of is single-component bicolored by observation 6.10, and no member of is component-wise bicolored by observation 6.10. Hence, no crucial -colorful graph can contain any member of as a colorful minor.
For the reverse direction let be a -colorful graph that does not contain any member of as a colorful minor. Due to lemma 6.15 and by lemmata 6.4 and 6.7 we know that is color-facial and color-segmented. Moreover, by observation 6.10 we have that is single-component bicolored and by observation 6.13 it is also component-wise bicolored. Hence, is crucial. ∎
@@
We believe that these numbers are surprisingly tame. Indeed, for some small values of we get and
@@
It is not hard to see that for any permutation of the -segregated grid realizing on its set of colors contains an half-integral packing of size of the -segregated grid which also realizes on its set of colors. See figure 7 for an example.
@@
The proof is divided into three parts spanning the following three subsections. First we show that every connected and crucial -colorful graph has the Erdős-Pósa property. Then we prove that this can be extended to cover also all disconnected crucial -colorful graphs, and finally, we show that -colorful graphs that are not crucial do not have the Erdős-Pósa property.
6.3 The positive part of theorem 6.19: The connected case
@@
Let us begin with a small observation. For there exist permutations such that there exists some for which the -segregated grid realizing is not a colorful minor of the -segregated grid realizing for any possible choice for See figure 2 for three such segregated grids in the case where However, for there exists only one segregated grid over all. The case is also special since there exist two cyclic permutations on however, it is easily seen that both permutations give rise to the same family of segregated grids up to isomorphism.999In general, for every positive integer there are different cyclic permutations without direction. Moreover, there exists a non-negative integer such that every -segregated grid is contained as a colorful minor in the -segregated grid corresponding to one of these permutations, for every non-negative integer
The following is a simple observation on the behavior of grids. See figure 8 for an illustration.
@@
Combining proposition 6.21 with observation 6.20 immediately yields that for large -segregated grids yield large packings for any crucial -colorful graph. Indeed, this observation can easily be extended as follows.
@@
The above property is implied by crucial graphs being single-component bicolored. Finally, we also need to observe how being component-wise bicolored implies a similar property. See figure 9 for an illustration.
@@
Observe that corollary 6.24 combined with theorem 1.6 directly implies the following.
@@
We first utilize lemma 6.26 to show that connected crucial colorful graphs have the Erdős-Pósa property. For this, recall that planar graphs have the Erdős-Pósa property for minors due to a result of Robertson and Seymour [76].
@@
For every non-negative integer every connected and crucial -colorful graph has the Erdős-Pósa property with gap where is the gap from proposition 6.27.
@@
For the assertion follows directly from the classic result of Robertson and Seymour, i.e. proposition 6.27. Hence, we may assume
@@
We first call upon corollary 6.25. This allows us to assume that as otherwise, corollary 6.25 would immediately provide us with a packing of size of within – and therefore within
Hence, by applying lemma 6.26 we now either find a packing of size of in or there is a set of size at most such that no component of carries all colors and contains as a colorful minor. However, by our choice of and as well as the assumption that is connected, the above implies that no restricted subgraph of may contain as a colorful minor. Moreover, since none of the colors in have any influence on whether is a colorful minor of or not, it follows that also does not contain as a colorful minor and our proof is complete. ∎
6.4 The positive part of theorem 6.19: The disconnected case
We continue with the proof of theorem 6.19, showing how we can reduce the case of general crucial -colorful graphs to theorem 6.28 where is connected.
@@
Base of the induction: The base follows in a straightforward manner from theorem 6.28. This is true since for the base we may consider the case where and is connected.
@@
Consider a -colorful graph and set for all Then, assuming and are both -colorful graphs and for any subgraph is a colorful minor of if and only if is a colorful minor of Observe that either there is a packing of in of size or by calling upon corollary 6.25, we may conclude with a set such that and every component is restricted. By lemma 6.26 we conclude with a set of size at most such that no connected component of contains as a colorful minor.
@@
6.5 Negative part of theorem 6.19
In this section we show that if is not crucial then it does not have the Erdős-Pósa property. This is the last remaining part for the proof of theorem 6.19.
@@
We begin with a simple but fundamental observation.
@@
Due to lemmata 6.30 and 6.17, it is enough to prove that none of the colorful graphs in has the Erdős-Pósa property. We first examine the case where is a connected -colorful graph from Clearly cannot be a member of as proposition 6.27 already covers all non-planar graphs.
@@
Assume to the contrary that contains two disjoint subgraphs and such that is a colorful minor of for both For let be the graph obtained by adding the vertex and joining it to all vertices of with a non-empty palette under Clearly both and are subgraphs of and their only common vertex is Notice that each contains one of as a minor. By proposition 6.31 this means that must contain a minor that is not embeddable in the projective plane which is absurd.
@@
We now claim that, in both cases, it is impossible for two color intervals, one of color and one of color to host the endpoints of and To see this we assume that there exists a where all of the only interact with two color intervals and construct an auxiliary graph as follows: In case the of we add a new vertex and make it adjacent to both endpoints of each path Moreover, we add a cycle along the boundary of (see the left side of figure 10). In the case the of we add a new vertex and make it adjacent to the endpoints of the colored by We then introduce a second new vertex and make it adjacent to the endpoints of the colored by (see the right side of figure 10). In both cases, the resulting graph has an embedding in the plane, by drawing precisely the new vertices in the complement of with respect to the plane. However this not possible as, in any case, the resulting graph, by construction, contains either or as a minor. We conclude that each occupies vertices of at least three color intervals. Therefore, as before, we can identify each of the, at most color intervals in the boundary of and see the graphs as hyperedges of arity drawn as open disks in the interior of a closed disk. Similar to previous claims, we may now observe that As before, this is a contradiction to the Erdős-Pósa property of
@@
Let be a -colorful graph that is the disjoint union of two (non-trivial) paths and whose endpoints are colored 1, 3, and 2, 4 respectively and whose internal vertices have empty palettes. Consider an embedding of in the closed disk with one crossing so that the colored vertices are drawn on the boundary of We next consider the -multiplication of in as indicated in figure 11. The resulting graph has a drawing in with crossings where every vertex carries at most one color. Notice also that all the crossing points are homotopic in
@@
The two directions of theorem 6.19 now follow from theorem 6.29 and theorem 6.32.
As a special case we investigate the Erdős-Pósa property for -colorful graphs that are rainbow, i.e. the palette of every vertex is As a consequence of theorem 6.19 we obtain the following.
@@
Case is obvious as this is precisely proposition 6.27.
For case recall that a graph is outerplanar if and only if it excludes both and as minos. Observe that if a graph contains or as a minor then the -colorful rainbow graph contains one of the -colorful graphs or from as a colorful minor. This, by theorem 6.19, it would not have the Erdős-Pósa property. For the inverse, because the graph of every -colorful graph from contains either or as a minor, for every -colorful graph that contains a member of as a minor it holds that must also contain or as a minor (see figure 5).
@@
theorem 6.19 and observation 6.16 imply that, for every is the colorful minor obstruction set of the class of -colorful graphs that have the Erdős Pósa property.
Given some we define the class as follows: contains the -grid without any colors, contains the -segregated grid and, for contains every -colorful graph that is the disjoint union of the set of all non-isomorphic -segregated grids with some common color (see figure 12). From observation 6.22 and observation 6.23 it follows that every crucial -colorful graph is a colorful minor of some graph in for some Notice that, while for as there are exactly way to choose the common color. Moreover it is direct to verify that every -colorful graph in is crucial. This permits us to restate theorem 6.19 as follows.

### v2

@@ 4.1 Colors in a surface
We should stress that the original formulation of proposition˜4.1 is not stated in terms of -colorful graphs and -colorful boundary indices. However, a variant like proposition˜4.1 is straightforward to obtain as in the original proof the boundary index is only used to assign a coloring to the non-vortex cells of a -rendition of which in this case would depend on Besides that, would not play any other role in its proof and can safely be disregarded thereafter.
@@ 4.2 Locally excluding a rainbow grid
Proposition˜4.1 allows us to immediately deduce a strong local structure theorem for -colorful graphs excluding rainbow grids.
@@
We proceed by applying proposition˜4.1 with and Notice that this means that This yields one of two outcomes as follows.
@@
Let be the minor model of and notice that is a truncation of We now apply theorem˜3.1. If this results in a colorful minor model of a rainbow we have found every -colorful rainbow graph on vertices as a colorful minor in and thus, in particular, we have found the rainbow -grid. Hence, we may assume that we enter the second outcome of theorem˜3.1 which yields a set of size at most
@@
4.3 The proof of Theorem˜1.4
We are now ready for the proof of Theorem˜1.4. The proof will largely follow a similar strategy as the one for theorem˜1.2. Indeed, we will prove a stronger version similar to theorem˜3.6 and its strengthened form theorem˜3.7. Theorem˜1.4 will then follow as a corollary. This slightly stronger version reads as follows:
@@
As announced, we strengthen theorem˜4.3 further to allow for a smooth induction.
@@
First, we define the functions involved as follows where is the constant from proposition˜2.3.
@@
We now proceed by induction on and start by discussing two fundamental cases. Exactly as in our proof of theorem˜3.7, we first discuss the two fundamental cases where either has only few vertices, or These two cases are fully identical to the ones in theorem˜3.7 and we therefore omit them here. Then two further cases follow:
Case 1, where has a small balanced separator, and Case 2, where does not have a small balanced separator. As before, for Case 1 there is no difference to the proof of theorem˜3.7 and thus, we may omit it here.
@@
Since we may now apply proposition˜2.3 to and obtain a -wall such that This further allows for the application of lemma˜4.2 to the wall in In the following we distinguish three cases based on the three possible outcomes of lemma˜4.2.
@@ 5 Excluding segregated grids
In this section, we are concerned with the proof of theorem˜1.3, which is an analogue of the Grid Theorem of Robertson and Seymour for torso treewidth. Similar to the organization of section˜4, we begin with some preliminary results and a “local” variant of the structure theorem before proceeding to the actual proof of theorem˜1.3.
@@ 5.1 A homogeneous flat wall
This allows us to call upon proposition˜2.4 to find either a model of controlled by in which case we would be done, or a set of size less than and an -subwall that is flat in
@@ 5.2 Locally excluding segregated grids
As it is custom, we begin by setting up our numbers. Let be the constant from proposition˜2.3 and hfw be the function from lemma˜5.1.
As a first step, we apply proposition˜2.3 to and obtain a -wall Next, we apply lemma˜5.1 to This either results in a -minor model with or in a set of size at most together with a homogeneous flat wall in of order
@@
We now apply theorem˜3.1. This either yields a rainbow as a colorful minor in or a set of size at most such that the -big component of is restricted. With being a truncation of the second outcome matches the second outcome of our assertion. Hence we may assume to find a rainbow as a colorful minor. This, however, means that contains every -colorful graph on vertices as a colorful minor and this includes all -segregated grids. Hence, we are done with this case.
@@
Now let for some brick of Since is homogeneous, we know that also is homogeneous, so the choice of here is arbitrary. Let and for each let us denote by the set We now apply lemma˜3.2 to the sets and the set asking for many paths for each member of This yields one of two outcomes. Either we find a linkage of order such that contains a --linkage of order for each and the paths in are all internalle vertex-disjoint from or there is a set of size less than such that there exists some and there is no --path in
@@
We claim that there exists an --linkage of size such that all internal vertices of the paths of belong to To see this, notice that in the absence of such a linkage proposition˜2.1 would imply the existence of a separation of order at most between and However, each vertex of belongs to a horizontal path of that can be extended in at least different ways, intersecting only in the horizontal path, to a path that intersects all cycles at least of Hence, such a separator must leave at least one cycle of one way to get from some vertex of to and one way to get from to intact. Hence, deleting any set of size at most still leaves at least one --path intact and therefore, and cannot be separated by fewer than vertices in
Let be the perimeter of and be the -subwall of that avoids It is now easy to see that together with the paths from the paths from the horizontal paths of and the bricks of that intersect we may construct a -segregated grid as a colorful minor of See figure˜4 for an illustration.
@@
5.3 A Grid Theorem for torso treewidth
We are now ready for the proof of theorem˜1.3. Similar to our proofs of Theorems˜3.6 and 1.4 we do this by proving a slightly different and stronger statement which, in turn, is implied by a further strengthening that lends itself better to induction.
@@ 5.3 A Grid Theorem for torso treewidth
contains a -segregated grid as a colorful minor and the torso treewidth of is at least or
@@
Notice that the tree-decomposition from the second outcome of theorem˜5.3 is a witness for the torso treewidth of being at most Hence, theorem˜5.3 does indeed act as an analogue of the Grid Theorem of Robertson and Seymour for torso treewidth.
We first prove that -segregated grids indeed act as lower bounds on the torso treewidth of a graph.
@@
For all positive integers and and every -colorful graph if contains a -segregated grid as a colorful minor, then the torso treewidth of is at least
@@
All that is left to do is to prove the following strengthening of the upper bound part of theorem˜5.3.
@@
We now proceed by induction on and start by discussing two fundamental cases. Exactly as in our proof of theorem˜3.7, we first resolve the two fundamental cases where either has only few vertices, or in precisely the same way as before. Then two further cases follow:
Case 1, where has a small balanced separator, and Case 2, where does not have a small balanced separator. As before, for Case 1 there is no difference to the proof of theorem˜3.7 and thus, we immediately move on to Case 2.
@@
In contrast to previous iterations of this proof, here we do not have to distinguish more subcases. We immediately may apply theorem˜5.2 to and which either yields a -segregated grid as a colorful minor in – and there would be nothing more to prove – or it provides us with a set with such that the -big component of is segregated.
@@ 6 The Erdős-Pósa property for -colorful graphs
With our structural theorems in place, we are now ready to proceed with the proof of theorem˜1.5. To this end, we first define, for every the sets and (section˜6.1); we then introduce the class of crucial colorful graphs (section˜6.1). The remainder of the section is devoted to proving that a -colorful graph has the Erdős-Pósa property if and only if it is crucial (theorem˜6.19). In the proof, the set serves as the (colorful minor) obstruction set for the class of all -colorful graphs that satisfy the Erdős-Pósa property. The “upper bound” part of the proof is presented in section˜6.3 (connected case) and section˜6.4 (non-connected case), while the “lower bound” is established in section˜6.5.
@@ 6.1 The sets and
We define by discarding all -colorful graphs with and every whose vertex set can be partitioned into two sets of size such that there are and For example, in the case where the only graph discarded is the one in the dashed square in figure˜5.
@@
If then See figure˜6 for an illustration.
@@
6.2 Crucial colorful graphs
@@
Recalling Observations˜6.2, 6.5, 6.8 and 6.11, one can see immediately that crucial colorful graphs are also closed under taking colorful minors.
@@
Taking into account lemma˜6.15, it is easy to verify the following.
@@
By observation˜6.16 we know that no member of contains another as a colorful minor. Moreover, by lemma˜6.4 we have that no member of is color-facial, no member of is color-segmented by observation˜6.7, no member of is single-component bicolored by observation˜6.10, and no member of is component-wise bicolored by observation˜6.10. Hence, no crucial -colorful graph can contain any member of as a colorful minor.
For the reverse direction let be a -colorful graph that does not contain any member of as a colorful minor. Due to lemma˜6.15 and by lemmata˜6.4 and 6.7 we know that is color-facial and color-segmented. Moreover, by observation˜6.10 we have that is single-component bicolored and by observation˜6.13 it is also component-wise bicolored. Hence, is crucial. ∎
@@
Interestingly, these numbers, being polynomially bounded, are rather tame. Indeed, for some small values of we get and
@@
It is not hard to see that for any permutation of the -segregated grid realizing on its set of colors contains an half-integral packing of size of the -segregated grid which also realizes on its set of colors. See figure˜7 for an example.
@@
The proof is divided into three parts spanning the following three subsections. First we show that every connected and crucial -colorful graph has the Erdős-Pósa property. Then we prove that this can be extended to cover also all non-connected crucial -colorful graphs, and finally, we show that -colorful graphs that are not crucial do not have the Erdős-Pósa property.
6.3 The positive part of theorem˜6.19: The connected case
@@
Let us begin with a small observation. For there exist permutations such that there exists some for which the -segregated grid realizing is not a colorful minor of the -segregated grid realizing for any possible choice for See figure˜2 for three such segregated grids in the case where However, for there exists only one segregated grid over all. The case is also special since there exist two cyclic permutations on however, it is easily seen that both permutations give rise to the same family of segregated grids up to isomorphism.101010In general, for every positive integer there are different cyclic permutations without direction. Moreover, there exists a non-negative integer such that every -segregated grid is contained as a colorful minor in the -segregated grid corresponding to one of these permutations, for every non-negative integer
The following is a simple observation on the behavior of grids. See figure˜8 for an illustration.
@@
Combining proposition˜6.21 with observation˜6.20 immediately yields that for large -segregated grids yield large packings for any crucial -colorful graph. Indeed, this observation can easily be extended as follows.
@@
The above property is implied by crucial graphs being single-component bicolored. Finally, we also need to observe how being component-wise bicolored implies a similar property. See figure˜9 for an illustration.
@@
Observe that corollary˜6.24 combined with theorem˜1.3 directly implies the following.
@@
We first utilize lemma˜6.26 to show that connected crucial colorful graphs have the Erdős-Pósa property. For this, recall that planar graphs have the Erdős-Pósa property for minors due to a result of Robertson and Seymour [76].
@@
For every non-negative integer every connected and crucial -colorful graph has the Erdős-Pósa property with gap where is the gap from proposition˜6.27.
@@
For the assertion follows directly from the classic result of Robertson and Seymour, i.e. proposition˜6.27. Hence, we may assume
@@
We first call upon corollary˜6.25. This allows us to assume that the torso treewidth of is at most as otherwise, corollary˜6.25 would immediately provide us with a packing of size of within – and therefore within
Hence, by applying lemma˜6.26 we now either find a packing of size of in or there is a set of size at most such that no component of carries all colors and contains as a colorful minor. However, by our choice of and as well as the assumption that is connected, the above implies that no restricted subgraph of may contain as a colorful minor. Moreover, since none of the colors in have any influence on whether is a colorful minor of or not, it follows that also does not contain as a colorful minor and our proof is complete. ∎
6.4 The positive part of theorem˜6.19: The non-connected case
We continue with the proof of theorem˜6.19, showing how we can reduce the case of general crucial -colorful graphs to theorem˜6.28 where is connected.
@@
Base of the induction: The base follows in a straightforward manner from theorem˜6.28. This is true since for the base we may consider the case where and is connected.
@@
Consider a -colorful graph and set for all Then, assuming and are both -colorful graphs and for any subgraph is a colorful minor of if and only if is a colorful minor of Observe that either there is a packing of in of size or by calling upon corollary˜6.25, we may conclude with a set such that and every component is restricted. By lemma˜6.26 we conclude with a set of size at most such that no connected component of contains as a colorful minor.
@@
6.5 Negative part of theorem˜6.19
In this section we show that if is not crucial then it does not have the Erdős-Pósa property. This is the last remaining part for the proof of theorem˜6.19.
@@
We begin with a simple but important observation.
@@
Due to lemmata˜6.30 and 6.17, it is enough to prove that none of the colorful graphs in has the Erdős-Pósa property. We first examine the case where is a connected -colorful graph from Clearly cannot be a member of as proposition˜6.27 already covers all non-planar graphs.
@@
Assume to the contrary that contains two disjoint subgraphs and such that is a colorful minor of for both For let be the graph obtained by adding the vertex and joining it to all vertices of with a non-empty palette under Clearly both and are subgraphs of and their only common vertex is Notice that each contains one of as a minor. By proposition˜6.31 this means that must contain a minor that is not embeddable in the projective plane which is absurd.
@@
We now claim that, in both cases, it is impossible for two color intervals, one of color and one of color to host the endpoints of and To see this we assume that there exists a where all of the only interact with two color intervals and construct an auxiliary graph as follows: In case the of we add a new vertex and make it adjacent to both endpoints of each path Moreover, we add a cycle along the boundary of (see the left side of figure˜10). In the case the of we add a new vertex and make it adjacent to the endpoints of the colored by We then introduce a second new vertex and make it adjacent to the endpoints of the colored by (see the right side of figure˜10). In both cases, the resulting graph has an embedding in the plane, by drawing precisely the new vertices in the complement of with respect to the plane. However this not possible as, in any case, the resulting graph, by construction, contains either or as a minor. We conclude that each occupies vertices of at least three color intervals. Therefore, as before, we can identify each of the, at most color intervals in the boundary of and see the graphs as hyperedges of arity drawn as open disks in the interior of a closed disk. Similar to previous claims, we may now observe that As before, this is a contradiction to the Erdős-Pósa property of
@@
Let be a -colorful graph that is the disjoint union of two (non-trivial) paths and whose endpoints are colored 1, 3, and 2, 4 respectively and whose internal vertices have empty palettes. Consider an embedding of in the closed disk with one crossing so that the colored vertices are drawn on the boundary of We next consider the -multiplication of in as indicated in figure˜11. The resulting graph has a drawing in with crossings where every vertex carries at most one color. Notice also that all the crossing points are homotopic in
@@
The two directions of theorem˜6.19 now follow from theorem˜6.29 and theorem˜6.32.
As a special case we investigate the Erdős-Pósa property for -colorful graphs that are rainbow, i.e. the palette of every vertex is As a consequence of theorem˜6.19 we obtain the following.
@@
Case is obvious as this is precisely proposition˜6.27.
For case recall that a graph is outerplanar if and only if it excludes both and as minos. Observe that if a graph contains or as a minor then the -colorful rainbow graph contains one of the -colorful graphs or from as a colorful minor. This, by theorem˜6.19, it would not have the Erdős-Pósa property. For the inverse, because the graph of every -colorful graph from contains either or as a minor, for every -colorful graph that contains a member of as a minor it holds that must also contain or as a minor (see figure˜5).
@@
theorem˜6.19 and observation˜6.16 imply that, for every is the colorful minor obstruction set of the class of -colorful graphs that have the Erdős Pósa property.
Given some we define the class as follows: contains the -grid without any colors, contains the -segregated grid and, for contains every -colorful graph that is the disjoint union of the set of all non-isomorphic -segregated grids with some common color (see figure˜12). From observation˜6.22 and observation˜6.23 it follows that every crucial -colorful graph is a colorful minor of some graph in for some Notice that, while for as there are exactly ways to choose the common color. Moreover it is direct to verify that every -colorful graph in is crucial. This permits us to restate theorem˜6.19 as follows.